### v4

# Contrastive Alignment with Semantic Gap-Aware Corrections in Text-Video Retrieval

Abstract
Recent advances in text-video retrieval have been largely driven by contrastive learning frameworks. However, existing methods overlook a key source of optimization tension: the separation between text and video distributions in the representation space—referred to as the modality gap—and the prevalence of false negatives in batch sampling. These factors lead to conflicting gradients under the InfoNCE loss, impeding stable alignment. To mitigate this, we propose GARE—a Gap-Aware Retrieval framework that introduces a learnable, pair-specific increment $\Delta_{ij}$ between text $t_{i}$ and video $v_{j}$ to offload the tension from the global anchor representation. We first derive the ideal form of $\Delta_{ij}$ via a coupling multivariate first-order Taylor approximation of the InfoNCE loss under a trust-region constraint, revealing it as a key mechanism for resolving gradient conflicts by guiding updates along a locally optimal descent direction in the coupled optimization landscape. Due to the expensive cost of directly approximate $\Delta_{ij}$, we introduce a lightweight neural module conditioned on the semantic gap between each video-text pair, enabling structure-aware correction guided by gradient supervision. To further stabilize learning and promote interpretability, we regularize $\Delta$ via three components: a trust-region constraint regularization to prevent oscillations, a directional diversity term to expand the semantic difference space, and an information bottleneck over $\Delta$ to restrict redundant information. Experiments across four retrieval benchmarks show that GARE consistently improves alignment accuracy and robustness to noisy supervision, confirming the effectiveness of gap-aware tension unloading. Code is available at https://github.com/musicman217/GARE-text-video-retrieval.
School of Computer Science and Information Engineering,
Hefei University of Technology, Hefei, China
{j.xiao_hfut, chenghao}@mail.hfut.edu.cnzjsonghfut@gmail.comzdszds534@gmail.com{zzhu, lijia}@hfut.edu.cnhongrc.hfut@gmail.com
1 Introduction
Text-video retrieval (TVR) [48] is a fundamental task in video understanding, aiming to retrieve relevant videos given a text query [29, 28, 39, 15]. With the proliferation of short video platforms, this task has attracted growing research interest. In recent years, vision-language pretraining models such as CLIP [33] have shown great success in cross-modal representation alignment, demonstrating strong performance on various retrieval benchmarks. These models typically learn a shared embedding space by aligning visual and textual modalities through large-scale contrastive learning [43, 18, 8, 9, 17, 40, 19, 27], and have thus become a popular backbone in TVR systems.
Despite the empirical success of contrastive learning in text-video retrieval, existing methods largely overlook two fundamental sources of optimization tension. First, the significantly modality gap between text and video representations: the two modalities often reside in disjoint regions of the embedding space [27, 47], with markedly different semantic structures and high distributional divergence (e.g., large KL divergence between modality-wise feature distributions). This separation creates an inherent geometric and semantic mismatch that cannot be easily bridged by global metric learning. Second, the prevalence of false negatives: semantically similar but unlabeled text-video pairs—further injects noisy gradient signals during training [10, 7], as they are incorrectly treated as hard negatives. These two factors jointly induce unstable or conflicting optimization forces during contrastive learning, especially under the coupled softmax structure of InfoNCE [32]. Without an explicit mechanism to account for these challenges, models often converge to suboptimal solutions, particularly on semantically ambiguous or structurally complex samples.
Figure 1: Tension and false negative challenge vs. our tension offloading strategy. (a) Due to modality gap [27] and false negatives [38], the gradients from negative samples partially overlap with the correct direction, leading to gradient tension around the anchor $t_{i}$. This restricts its optimization freedom and causes erroneous updates. (b) We offload part of the optimization pressure from $t_{i}$ to the learned increment $\Delta_{ij}$ by relaxing the gradient. This relieves gradient tension and reduces the impact of false negatives on $t_{i}$, as errors are partially absorbed by $\Delta$. Each $\Delta_{ij}$ captures a semantically meaningful correction of the text-video gap.
Figure 1(a) provides an overview of the optimization tension caused by modality gaps, where text and video embeddings lie in well-separated regions of the representation space—this repulsion tends to push $t_{i}$ away from the entire video manifold. As a result, the positive and negative terms impose a conflicting descent direction, both acting along the same region but with opposing effects. This constitutes a core source of optimization tension in contrastive learning. In addition, false negatives—semantically relevant but unlabeled pairs—introduce noisy repulsion that further distorts the optimization trajectory, especially in early training.
To address both issues, we introduce a pair-specific increment $\Delta_{ij}$ that acts as a learnable adjustment between $t_{i}$ and $v_{j}$. As shown in Figure 1(b), instead of enforcing all semantic correction on the anchor embedding $t_{i}$, we allow $\Delta_{ij}$ to absorb part of the optimization signal at the pair level. This design serves two complementary purposes: 1) it unloads the global optimization tension caused by the modality gap by decoupling the descent direction across different video-text pairs; 2) it buffers the local gradient noise introduced by false negatives, preventing anchor representations from drifting due to misaligned repulsion.
To ensure that $\Delta_{ij}$ contributes constructively to the loss landscape, we derive its ideal form via a first-order Taylor expansion of the contrastive loss under a trust-region constraint. While the exact solution is intractable due to inter-pair coupling, we approximate it using a lightweight neural module that conditions on the semantic gap $v_{j}-t_{i}$, allowing the model to proactively infer alignment adjustments before receiving backpropagated signals. This results in a closed-loop training dynamic: the model first estimates a gap-aware correction based on its prior belief, then updates this belief through feedback from contrastive supervision. Together, this enables more stable training and more accurate modeling of semantic structure under noise.
We further analyze the geometric behavior of the learned increment $\Delta_{ij}$. Empirically, we observe that $\Delta_{ij}$ tends to form an obtuse angle with the initial gap vector $v_{j}-t_{i}$, averaging around 1.6 radians. This indicates that the model is not merely copying the raw modality difference, but is instead learning a structure-aware correction that incorporates higher-level alignment priors. Moreover, we find that the Euclidean distance between paired embeddings increases after applying $\Delta_{ij}$, suggesting that the model leverages pair-level adjustment to release optimization tension and enable finer-grained control over semantic positioning on the unit sphere. Our contributions are as follows:
  • We analyze the gradient structure of the InfoNCE loss and reveal its inherent multi-variable coupling behavior via introduced multiple increments $\Delta$. By performing a first-order Taylor expansion under a local trust region, we derive an optimal update rule for each pairwise increment $\Delta_{ij}$ that aligns with the descent direction of InfoNCE.
  • Building on this formulation, we propose a gap-aware retrieval framework where each increment $\Delta_{ij}$ is predicted by a learnable network and directly participates in the forward pass. This design enables explicit offloading of gradient tension from text anchors, thereby mitigating misaligned updates caused by modality gap and false negatives.
  • We validate our method on four cross-modal retrieval benchmarks, i.e., MSR-VTT [45], DiDeMo [2], ActivityNet Captions [24], MSVD [6], demonstrating consistent improvements. Further analysis shows that the learned increments are semantically meaningful and geometrically structured, highlighting their role in stable and interpretable optimization.
2 Related work
Contrastive Learning and Modality Gap
Contrastive learning has become a foundational paradigm in multimodal representation learning. Wang and Isola [40] formalize contrastive learning via the principles of alignment and uniformity on the hypersphere, offering a geometric perspective on representation quality. Wang and Liu [38] show that contrastive loss is hardness-aware and temperature-sensitive, and reveal a trade-off between representation uniformity and semantic tolerance, highlighting the need to preserve meaningful structure among semantically similar samples. Liang et al.[27] investigate the modality gap in multimodal contrastive learning and attribute it to initialization imbalance and cone effects. False negatives have also been recognized as a key challenge in contrastive learning [10, 7], with solutions ranging from reweighting and elimination to dynamic detection and correction.
Text-Video Retrieval
Text-video retrieval is one of the prominent tasks  [13, 23, 37, 3, 16, 44, 31, 4, 21, 25, 41] in cross-modal learning. The majority of existing research  [26, 14, 22, 30, 42, 5, 46] in this area utilizes a mapping technique that aligns both text and video inputs within a shared latent space to facilitate direct similarity assessment. CLIP4Clip [29] is the first to adapt CLIP [33] for video-text retrieval via temporal frame aggregation. TS2-Net [28] improves temporal modeling through token shift and selection. X-Pool [15] introduces text-guided pooling to highlight salient video tokens. HBI [19] values possible correspondences between frames and words using Banzhaf interaction for sensitive and explainable cross-modal alignment. EMCL [19] introduces an expectation-maximization [11] framework to learn a compact latent space where video and text features are represented as linear combinations of shared bases. This decomposition reduces the rank of the latent space, mitigating the modality gap and enhancing semantic alignment. Unlike prior works that refine matching structures, our method analyzes the gradient form of InfoNCE and introduces a learnable gap-aware increment $\Delta_{ij}$ to offload optimization tension, enabling structured optimization in a trust-region-aware formulation.
3 Approach
We propose a gap-aware retrieval framework that introduces a learnable, pair-specific increment $\Delta_{ij}$ to adjust text features toward video embeddings or vice versa. To alleviate optimization tension and false negatives, we derive $\Delta_{ij}$ via a first-order Taylor expansion of the InfoNCE loss and approximate it using a neural module $\psi$ conditioned on the semantic gap. We further regularize $\Delta_{ij}$ in norm, direction, and distribution to enhance its structure and interpretability.
3.1 Preliminaries
Task Definition
Given a dataset of $N$ paired video-text examples $\{(v_{i},t_{i})\}_{i=1}^{N}$, the goal of text-video retrieval is to learn a pair of encoders: a visual encoder $\phi_{v}(\cdot)$ and a text encoder $\phi_{t}(\cdot)$, that map inputs into a shared embedding space. The similarity between any pair $(v_{j},t_{i})$ is computed using a scoring function, typically the cosine similarity:
$s_{ij}=\frac{\phi_{t}(t_{i})^{\top}\phi_{v}(v_{j})}{\|\phi_{t}(t_{i})\|\cdot\|\phi_{v}(v_{j})\|},$ (1)
where $\phi_{t}(t_{i})\in\mathbb{R}^{D}$ and $D$ is dimension. During training, contrastive learning is applied to increase the similarity of matched pairs while decreasing that of mismatched ones. At inference, retrieval is performed by ranking all candidate texts (or videos) for a given query video (or text) based on similarity scores. For brevity, we henceforth refer to $\phi_{t}(t)$ as $t$, $\phi_{v}(v)$ as $v$.
Optimization Tension and First-order Increment Modeling
As discussed in the introduction, contrastive optimization in video-text retrieval suffers from gradient tension caused by modality gap and noise from false negatives. These factors hinder stable optimization of the anchor representation $t_{i}$, which must simultaneously align with the positive $v_{i}$ and repel all negatives $v_{j\neq i}$. Formally, the per-anchor InfoNCE loss is given by:
$\mathcal{L}_{i}=-\log\frac{e^{\cos(t_{i},v_{i})/\tau}}{\sum_{j=1}^{B}e^{\cos(t_{i},v_{j})/\tau}},$ (2)
and its gradient with respect to $t_{i}$ is:
$\nabla_{t_{i}}\mathcal{L}_{i}=\sum_{j=1}^{B}(p_{ij}-y_{ij})\cdot\left[\frac{v_{j}}{\|t_{i}\|\|v_{j}\|}-\cos(t_{i},v_{j})\cdot\frac{t_{i}}{\|t_{i}\|^{2}}\right],$ (3)
where $p_{ij}=\text{softmax}(\cos(t_{i},v_{j})/\tau)$ and $y_{ij}=\mathds{1}_{[j=i]}$. As observed in prior works such as [38], the magnitudes of the positive gradient coefficient $|p_{ii}-1|$ and the aggregate negative terms $\sum_{j\neq i}p_{ij}$ are comparable. In cases of strong modality gap, the attraction from the positive and the repulsion from negatives tend to pull in opposing directions, creating intrinsic gradient tension on $t_{i}$.
To alleviate both challenges mentioned above, we introduce a pair-specific increment $\Delta_{ij}$ that enables the model to locally adjust the relative positioning between each text-video pair. This increment serves a dual role: it offloads optimization tension by distributing alignment corrections across pairs, and buffers gradient noise by absorbing false-negative interference before it propagates into the anchor embedding. For clarity, we first consider the case where $\Delta_{ij}$ is applied to the text side, yielding an adjusted representation:
$t_{\Delta_{ij}}=t_{i}+\Delta_{ij}.$ (4)
This formulation allows for certain perturbations at the given anchor $t_{i}$ corresponding to different $v_{j}$ respectively. In practice, depending on dataset-specific modality characteristics, the same mechanism can also be applied to the video side (e.g., $v_{j}+\Delta_{ij}$). However, this flexibility raises a central question: how should each $\Delta_{ij}$ be optimized to effectively reduce the contrastive loss?
A naive approach would linearize the loss around the original anchor $t_{i}$ (i.e., $\Delta_{i*}=0$), treating each $\Delta_{ij}$ as an independent perturbation. However, under this formulation, the gradient $\nabla_{\Delta_{ij}}\mathcal{L}_{i}$ depends only on the static similarity $\cos(t_{i},v_{j})$, failing to account for the influence of other $\Delta_{ik}$ in the same batch. This leads to decoupled gradients that ignore the softmax coupling intrinsic to InfoNCE. More importantly, expanding the loss only at $t_{i}$ results in a univariate approximation, which does not reflect how different $\Delta_{ij}$ collectively reshape the similarity ranking across all $v_{j}$. Since our goal is to adjust the relative structure of the entire pair set for each anchor, we require a multivariate formulation where all $\Delta_{ij}$ are coupled and optimized under a shared comparison scale.
To address this, we treat all $\Delta_{ij}$ as jointly optimized variables, and reinterpret the per-anchor contrastive loss $\mathcal{L}_{i}$ as multivariate function over the full set $\{\Delta_{ij}\}_{j=1}^{B}$:
$\mathcal{L}_{i}(\Delta_{i1},\dots,\Delta_{iB})=-\log\frac{e^{s_{ii}/\tau}}{\sum_{k=1}^{B}e^{s_{ik}/\tau}},\quad s_{ij}=\cos(t_{i}+\Delta_{ij},v_{j}).$ (5)
Unlike standard per-sample optimization, the softmax structure couples all $\Delta_{ij}$, meaning the gradient of any single increment depends on the values of the others. To capture this dependency, we perform a multivariate first-order Taylor expansion around a prior coupled state $\Delta_{i*}^{(0)}=\{\Delta_{ij}^{(0)}\}_{j=1}^{B}$, where all increments are nonzero. In practice, we treat this prior as the current state at iteration $t$ — i.e., we let $\Delta_{i*}^{(t)}:=\Delta_{i*}^{(0)}$ — and analyze the local descent behavior from this reference point. This converts our optimization into an iterative process, where each step refines the current set of coupled increments. The resulting linear approximation is:
$\mathcal{L}_{i}(\Delta_{i*})\approx\mathcal{L}_{i}(\Delta_{i*}^{(t)})+\sum_{j=1}^{B}\left[\nabla_{\Delta_{ij}}\mathcal{L}_{i}(\Delta_{i*}^{(t)})\right]^{\top}(\Delta_{ij}-\Delta_{ij}^{(t)}).$ (6)
Crucially, this expansion preserves the softmax-induced coupling: each gradient term $\nabla_{\Delta_{ij}}\mathcal{L}_{i}$ is evaluated assuming that all other increments $\Delta_{ik\neq j}$ remain fixed at their nonzero states. The resulting linearized loss defines a local descent landscape for each $\Delta_{ij}$. To ensure reliable updates within this approximation, we impose a trust-region constraint that bounds the update magnitude:
$\|\Delta_{ij}\|\leq\varepsilon,$ (7)
which reflects our belief that meaningful corrections should occur within a localized region around $t_{i}$. Under this constraint, the optimal update direction for each $\Delta_{ij}$ corresponds to steepest descent along its local gradient:
$\Delta_{ij}^{*(t+1)}=\Delta_{ij}^{(t)}-\alpha_{ij}^{(t)}\cdot\frac{\nabla_{\Delta_{ij}}\mathcal{L}_{i}(\Delta_{i*}^{(t)})}{\|\nabla_{\Delta_{ij}}\mathcal{L}_{i}(\Delta_{i*}^{(t)})\|},$ (8)
where step size $\alpha_{ij}^{(t)}$ is analytically determined to ensure $\|\Delta_{ij}^{(t+1)}\|\leq\varepsilon$ and derived via the Cauchy–Schwarz inequality to project onto the trust-region boundary (see Appendix D for details). This yields the steepest descent direction of the linearized loss, evaluated at a coupled batch state $\Delta_{i*}^{(t)}$, where all increments are fixed but nonzero. While conceptually similar to coordinate-wise proximal optimization, directly applying this solution is impractical due to two key challenges: 1) gradient coupling: each $\nabla_{\Delta_{ij}}\mathcal{L}_{i}$ depends on all other $\Delta_{ik}$, making updates interdependent and unstable; 2) scale ambiguity: the optimal trust-region radius $\varepsilon$ may vary across pairs depending on semantic difficulty or training stage. To address this, we adopt a learnable network that directly predicts the coupled increment state $\Delta_{i*}^{(t)}$ based on the semantic gap $v_{j}-t_{i}$. The subsequent update to $\Delta_{i*}^{(t+1)}$ is implicitly performed via backpropagation on the InfoNCE loss. Since the gradient of the loss naturally aligns with the steepest descent direction, this formulation allows us to transform the iterative update rule into a trainable network prediction problem, enabling amortized, structure-aware optimization over the full batch of pairwise increments.
3.2 Gap-Aware Increment Modeling via Pair-Specific $\Delta_{ij}$
To make this optimization tractable, we replace the explicit iterative updates with a learnable function $\psi$ that directly predicts each increment $\Delta_{ij}^{(t)}$. Specifically, we amortize the descent process by training a parameterized network $\psi$ to generate the current coupled increment state based on a pairwise semantic difference and a modality-dependent context:
$\Delta_{ij}^{(t)}=\psi\left(\eta(v_{j}-t_{i}),c;\Theta^{(t)}\right),$ (9)
where $\eta\in\{-1,+1\}$ determines the direction of the semantic gap depending on which modality receives the increment. The function $\psi$ is implemented as a single-layer cross-attention module, where the semantic difference $\eta(v_{j}-t_{i})$ serves as the query, and the context vector $c$—defined based on dataset characteristics—is used as both the key and the value. This allows $\psi$ to generate $\Delta_{ij}$ in a modality- and structure-aware manner, conditioned on the local pairwise discrepancy and auxiliary context. Rather than computing explicit gradient steps, $\psi$ learns to output increments that implicitly reflect the local descent direction. Since $\Delta_{ij}$ is embedded directly in the InfoNCE loss, the network is trained end-to-end via backpropagation, allowing it to align its output with the true coupled gradient field. As a result, $\psi$ captures both structure-aware priors and semantic context, producing informed corrections that reduce optimization tension and improve alignment robustness across diverse retrieval scenarios. We now consider the case where $c=v_{j}$ and the increment $\Delta_{ij}$ is added to the text side. In this setting, the two variables $t_{i}$ and $\Delta_{ij}$ share the same gradient flow:
$\nabla_{t_{i}}\mathcal{L}_{i}=\sum_{j=1}^{B}\nabla_{t_{\Delta_{ij}}}\mathcal{L}_{i},\qquad\nabla_{\Delta_{ij}}\mathcal{L}_{i}=\nabla_{t_{\Delta_{ij}}}\mathcal{L}_{i},$ (10)
where the gradient with respect to each perturbed anchor is given by:
$\nabla_{t_{\Delta_{ij}}}\mathcal{L}_{i}=(p_{ij}-y_{ij})\cdot\left[\frac{v_{j}}{\|t_{\Delta_{ij}}\|\cdot\|v_{j}\|}-\cos(t_{\Delta_{ij}},v_{j})\cdot\frac{t_{\Delta_{ij}}}{\|t_{\Delta_{ij}}\|^{2}}\right].$ (11)
This formulation introduces a gradient redistribution effect: unlike standard InfoNCE where all gradients act directly on the shared anchor $t_{i}$, our pair-specific design allows each negative video $v_{j}$ to influence only its associated increment $\Delta_{ij}$. The positive pair $(t_{i},v_{i})$ contributes attraction gradients to both $t_{i}$ and $\Delta_{ii}$, facilitating alignment; meanwhile, each negative pair $(t_{i},v_{j}),j\neq i$, applies repulsion primarily to $\Delta_{ij}$. This leads to two key benefits: 1) tension relief — $\Delta_{ii}$ can absorb the alignment shift toward $v_{i}$, reducing the burden on $t_{i}$ to resolve large modality gaps in a single step; 2) false-negative suppression — repulsion from semantically similar negatives is redirected into their respective increments, preserving the semantic stability of the anchor representation. We apply this strategy under a symmetric InfoNCE loss:
$\mathcal{L}_{\text{info}}=-\frac{1}{2}\left(\frac{1}{B}\sum_{i=1}^{B}\log\frac{e^{s_{ii}/\tau}}{\sum_{j=1}^{B}e^{s_{ij}/\tau}}+\frac{1}{B}\sum_{j=1}^{B}\log\frac{e^{s_{ii}/\tau}}{\sum_{i=1}^{B}e^{s_{ij}/\tau}}\right),$ (12)
This loss function maximizes the similarity of positive pairs $s(t_{i}+\Delta_{ii},v_{i})$ and minimizes the similarity of negative pairs.
Norm-Based Regularization of Trust-Region Radii
Each pair-specific increment $\Delta_{ij}$, predicted by the network $\psi$, approximates the optimal update direction under the first-order Taylor approximation. Its norm $\varepsilon_{ij}=\|\Delta_{ij}\|_{2}$ thus serves as a learned trust-region radius, regulating the extent of deviation that $t_{i}$ can take—regardless of direction—in order to absorb local alignment tension and semantic adjustment. To encourage structure in these radii, we regularize the intra-anchor variance of $\varepsilon_{ij}$ across each text anchor. Intuitively, semantically close pairs should have smaller $\varepsilon_{ij}$, while hard negatives benefit from larger ones—yet such structure must be inferred implicitly. We therefore promote norm diversity by minimizing the expected negative variance across anchors:
$\mathcal{L}_{\varepsilon}=\mathbb{E}_{t_{i}\sim\mathcal{B}_{t}}\left[-\text{Var}\left(\{\varepsilon_{ij}\}_{j=1}^{B}\right)\right],$ (13)
where $\mathcal{B}_{t}$ denotes the batch of text anchors. This encourages sharper separation between low-confidence and high-confidence pairs, without relying on external supervision. To prevent instability, we apply a lower bound:
$\mathcal{L}_{\varepsilon}=\max(\mathcal{L}_{\varepsilon},-\lambda),\quad\lambda>0.$ (14)
This simple yet effective regularization sharpens the implicit trust-region structure learned by $\psi$, guiding $\Delta_{ij}$ to reflect per-pair semantic variability while maintaining stable training dynamics. In the following section, we further complement this with a directional diversity loss to explicitly enhance the angular resolution of $\Delta_{ij}$ under each anchor.
Directional Diversity Regularization
To enhance the expressiveness of the learned increments $\Delta_{ij}$, we introduce a directional regularization that encourages the directions of $\{\Delta_{ij}\}_{j=1}^{B}$ under each anchor $t_{i}$ to be diverse. This helps the model assign distinct update directions to different candidate videos, improving representational resolution and mitigating mode collapse.
We normalize each increment to obtain unit vectors $z_{ij}=\frac{\Delta_{ij}}{\|\Delta_{ij}\|_{2}}$ and define the regularization loss as the expected angular similarity across all anchor-specific direction sets:
$\mathcal{L}_{\text{dir}}=\mathbb{E}_{t_{i}\sim\mathcal{B}_{t}}\left[\log\mathbb{E}_{j,k}\left[\exp\left(-\sigma\cdot(1-\langle z_{ij},z_{ik}\rangle)\right)\right]\right],$ (15)
where $\sigma$ is a scale factor to control uniformity. This loss softly penalizes directional concentration, while still allowing nearby directions for semantically similar negatives—preserving flexibility under uncertainty. Combined with the norm-based regularization in Section 3.2.1, this term enables fine-grained control over both the magnitude and direction of each increment $\Delta_{ij}$, leading to more stable and structure-aware optimization.
KL Regularization via Information Bottleneck
To regularize the distribution of increments $\Delta_{ij}$ across text inputs for each video anchor $v_{j}$, we introduce a Information Bottleneck (IB) [1, 35] via a KL divergence loss that aligns the empirical distribution with a standard normal prior:
$\mathcal{L}_{\text{IB}}=\mathbb{E}_{v_{j}\sim\mathcal{B}_{v}}\left[\mathrm{KL}\left(\mathcal{N}(\mu_{j},\sigma_{j}^{2})\,\|\,\mathcal{N}(0,I)\right)\right],$ (16)
where $\mu_{j}$ and $\sigma_{j}^{2}$ are the mean and variance of $\{\Delta_{ij}\}_{i=1}^{B}$ over the text batch for each $v_{j}$. This loss can be interpreted through the lens of the Information Bottleneck principle: it encourages $\Delta$ to retain minimal information about $v_{j}$ beyond what is necessary for pairwise alignment. By bounding the mutual information $I(\Delta;v_{j})$, the model avoids encoding redundant video-specific patterns in $\Delta$. This regularization complements the norm and direction constraints by promoting centered, isotropic, and information-efficient corrections, further enhancing generalization and stability.
Although the input configuration of $\psi$ may vary across datasets, we consistently apply the KL regularization along the video dimension, as video-side representations typically exhibit higher redundancy and local similarity. Regularizing the set of increments associated with each video anchor helps suppress noise and stabilizes learning by enforcing consistent structure across semantically similar pairs. See Appendix.B for an ablation study.
4 Experiment
Datasets, Metrics and Implementation Details
Datasets. We evaluate our method on four standard text-video retrieval benchmarks: MSR-VTT [45], DiDeMo [2], MSVD [6], and ActivityNet Captions [24]. MSR-VTT contains 10K videos with 20 captions each; we follow the 1K-A test split. DiDeMo includes 10K videos segmented into 5-second clips, each annotated with multiple sentences. MSVD consists of 1.9K short video clips with English captions. ActivityNet Captions provides dense annotations for 20K long-form videos with multiple temporally grounded descriptions.
Metrics. We choose Recall at rank K={1, 5, 10} (R@K), Median Rank (MdR), and mean rank (MnR) to evaluate the retrieval performance.
Implementation Details. We adopt CLIP (ViT-B/32) [33] as the base bi-encoder, equipped with a 4-layer Temporal Transformer [36] for video encoding. Following prior works [29, 15, 26, 39], we use 32-word captions and 12 video frames for MSR-VTT and MSVD, and 64-word captions with 64 frames for DiDeMo and ActivityNet Captions due to their longer video durations. We use the Adam optimizer [12] with linear warm-up, as in prior works. The learning rate is set to $1\text{e}{-7}$ for CLIP’s text and visual encoders, and $1\text{e}{-4}$ for all other modules (except $3\text{e}{-4}$ for ActivityNet). We set $\tau=0.01$, $\sigma=2$, and $\lambda=0.5$. All experiments use a batch size of 128. We train the model for 5 epochs on MSR-VTT, MSVD, and DiDeMo, and 10 epochs on ActivityNet Captions. All experiments are conducted on 4 to 8 GPUs including RTX 4090, A100 and V100.
Table 1: Comparison results on MSRVTT dataset on Text-to-Video Retrieval and Video-to-Text Retrieval.
Methods Text-to-Video Retrieval Video-to-Text Retrieval
R@1↑ R@5↑ R@10↑ MdR↓ MnR↓ R@1↑ R@5↑ R@10↑ MdR↓ MnR↓
CLIP4Clip [29] 44.5 71.4 81.6 2.0 15.3 42.7 70.9 80.6 2.0 11.6
X-Pool [15] 46.9 72.8 82.2 2.0 14.3 44.4 73.3 84.0 2.0 9.0
TS2-Net [28] 47.0 74.5 83.8 2.0 13.0 45.3 74.1 83.7 2.0 9.2
EMCL-Net [19] 46.8 73.1 83.1 2.0 12.8 46.5 73.5 83.5 2.0 8.8
UATVR [14] 47.5 73.9 83.5 2.0 12.3 46.9 73.8 83.8 2.0 8.6
ProST [26] 48.2 74.6 83.4 2.0 12.4 46.3 74.2 83.2 2.0 8.7
HBI [20] 48.6 74.6 83.4 2.0 12.0 46.8 74.3 84.3 2.0 8.9
Diffusion [22] 49.0 75.2 82.7 2.0 12.1 47.7 73.8 84.5 2.0 8.8
EERCF [34] 47.8 74.1 84.1 - - 44.7 74.2 83.9 - -
MPT [49] 48.3 72.0 81.7 - 14.9 46.5 74.1 82.6 - 11.8
Baseline 46.6 73.4 82.2 2.0 12.6 45.6 73.4 82.4 2.0 9.6
GARE (Ours) 49.1 74.7 83.6 2.0 12.0 48.6 75.3 85.3 2.0 8.5
Table 2: Comparison results on DiDeMo, ActivityNet Captions, and MSVD datasets on Text-to-Video Retrieval.
DiDeMo
Methods R@1 R@5 R@10 MnR
TS2-Net 41.8 71.6 82.0 14.8
CLIP4Clip 42.8 68.5 79.2 18.9
Diffusion 46.7 74.7 82.7 14.3
HBI 46.9 74.9 82.7 12.1
Baseline 45.4 74.3 82.0 12.3
GARE (Ours) 47.6 75.4 83.1 12.0
ActivityNet Captions
Methods R@1 R@5 R@10 MnR
CLIP4Clip 40.5 72.4 83.6 7.5
TS2-Net 41.0 73.6 84.5 8.4
MPT 41.4 70.9 82.9 7.8
HBI 42.2 73.0 84.6 6.6
Baseline 40.2 72.5 83.6 7.5
GARE (Ours) 42.6 73.2 84.8 6.6
MSVD
Methods R@1 R@5 R@10 MnR
CLIP4Clip 45.2 75.5 84.3 10.3
EMCL-Net 42.1 71.3 81.1 17.6
UATVR 46.0 76.3 85.1 10.4
Diffusion 46.6 75.9 84.1 15.7
Baseline 45.0 75.5 84.5 10.7
GARE (Ours) 46.4 76.1 84.5 10.6
Comparison with Other Methods
Table 1 and Table 2 shows the performance of our method across four standard text-video retrieval benchmarks. As seen, our approach consistently outperforms recent state-of-the-art methods on MST-VTT, ActivityNet and DiDeMo, and a comparable results in MSVD.
Ablative Analysis
Table 3: Ablation Studies on Modality Choice and Loss Combinations. left: We perform Text-to-Video Task on three datasets on different context modality type of our module $\psi$. right: We perform Text-to-Video Task on MSR-VTT for all losses ablation. First row refers to our baseline.
Context Modality Choice of $\psi$
Modality Dataset R@1↑ R@5↑ R@10↑ MnR↓
$c=t_{i}$ MSRVTT 47.4 73.5 82.1 12.9
$c=v_{j}$ MSRVTT 49.1 73.3 82.2 12.4
$c=t_{i}$ ActivityNet 42.6 73.6 84.4 6.8
$c=v_{j}$ ActivityNet 40.2 72.2 83.6 8.1
$c=t_{i}$ DiDeMo 46.5 74.3 82.6 12.3
$c=v_{j}$ DiDeMo 47.6 75.4 83.1 12.0
Losses Ablation
$\Delta$ $\mathcal{L}_{\textbf{IB}}$ $\mathcal{L}_{\varepsilon}$ $\mathcal{L}_{\textbf{dir}}$ R@1↑ R@5↑ R@10↑ MnR↓
Baseline 46.6 73.4 82.2 12.6
47.4 73.8 82.8 12.4
47.2 73.3 82.2 12.4
47.0 73.1 82.3 12.6
47.4 73.7 82.8 12.3
48.3 74.2 83.2 12.4
49.1 74.7 83.6 12.0
Losses design. We conduct ablation studies on the MSR-VTT 1k-A test set to assess the effectiveness of the proposed increment $\Delta$ and its associated regularizers. As shown in the right of Table 3, directly injecting $\Delta$ into the InfoNCE flow improves performance from 46.6 to 47.4, validating the benefit of gradient tension release via pairwise adjustment. Introducing the information bottleneck (IB) loss further boosts performance to 48.3, highlighting its role in guiding $\Delta$ toward semantically meaningful corrections. However, adding the norm constraint or directional diversity loss alone yields no gain. Notably, combining IB with both the norm and diversity regularizers leads to a substantial improvement, reaching 49.1. This suggests that controlling the update scale (via trust-region) and promoting directional diversity must work together to fully exploit the potential of $\Delta$. This aligns with our design intuition: the trust-region constraint controls the update scale (via $\varepsilon$) but does not regulate directionality, which can lead to redundant or entangled corrections. The diversity regularization ensures that increments associated with the same anchor are directionally distinct, encouraging more uniform and structured behavior around each anchor.
Context Modality Choice in $\psi$. We evaluate the effect of applying $\Delta_{ij}$ to either modality across datasets with contrasting characteristics. As shown in the left of Table 3, on MSR-VTT, injecting $\Delta$ on the text side ($t_{i}+\Delta_{ij}$) achieves the best performance, while applying it to the video side degrades results (R@1 drops to 47.4). This aligns with the fact that MSR-VTT contains many visually similar short clips, making it more effective to adjust the abstract text representation for fine-grained distinctions. Conversely, on ActivityNet, applying $\Delta$ to the video side leads to a notable performance boost, whereas modifying text harms results (R@1 $\approx$ 40). This is likely because the long but redundant videos are paired with rich, structured captions—making video-side adaptation more beneficial. These trends highlight the importance of aligning $\psi$’s modality choice with dataset structure. For DiDeMo, given its moderate video length (30s), we treat video as the anchor and set $c=v_{j}$ accordingly.
More ablation analysis including the lower bound $\lambda$ of $\mathcal{L}_{\varepsilon}$, the combination between context modality type $c$ of $\psi$ and $\eta$, the scale factor $\sigma$ of $\mathcal{L}_{\text{dir}}$ and the choice of anchor of $\mathcal{L}_{\text{IB}}$ can be found in Appendix.B.
5 Qualitative Analysis
To understand the effect of the learned increment $\Delta$, we conduct a qualitative analysis on the MSR-VTT 1k-A validation set, focusing on its impact on representation geometry and alignment behavior at inference. As shown in Figure 2(a), our method GARE yields lower cosine similarities between positive pairs than the baseline, indicating larger angular separation and improved uniformity on the unit hypersphere [40]. Figures 2(b) and 2(c) further show that while $t_{i}$ is concentrated around norm 12, the offset $t_{\Delta_{ij}}$ exhibits greater and more diverse magnitudes. This supports our view that $\Delta$ serves as a tension offloading mechanism, allowing $t_{i}$ to remain a stable prototype while $t_{\Delta_{ij}}$ absorbs pair-specific adjustments—particularly useful under gradient noise from false negatives. As shown in Figure 2(d), $t_{\Delta_{ij}}$ is generally farther from $v_{j}$ than $t_{i}$, indicating that the model does not simply pull pairs closer but expands the representation scale to allow more flexible alignment. This explains why, despite lower cosine similarity, GARE achieves better uniformity and geometric balance. Additional training-time analysis is provided in Appendix.C.
(a) Cosine similarity distribution of positive pairs.
(b) Norm distribution on negative pairs.
(c) Norms distribution on positive pairs.
(d) Distance shift: $t_{\Delta}$ vs. $t$.
Figure 2: Qualitative analysis on the MSR-VTT 1k-A validation set. $t_{\text{new}}$ denotes $t_{\Delta}$. Our method induces greater angular separation between positive pairs (a), redistributes $t_{\Delta}$ norms to release gradient tension (b, c), and pushes $t_{\Delta}$ outward from $v_{j}$ (d), promoting uniformity.
6 Conclusion
This paper investigates the optimization behavior of contrastive learning in text-video retrieval, focusing on the tension caused by modality gap and the noise introduced by false negatives. We show that such tension arises when gradients from negative pairs conflict with positive updates, restricting the optimization freedom of text anchors. To address this, we perform a first-order Taylor expansion of the InfoNCE loss and derive a pair-specific increment $\Delta_{ij}$ as an optimal adjustment direction within a trust region. By learning $\Delta_{ij}$ through a neural module conditioned on the semantic gap, and regularizing its structure, we enable more stable and interpretable optimization. Experiments on standard retrieval benchmarks confirm the effectiveness of this approach, and qualitative results demonstrate that $\Delta$ captures meaningful geometry for resolving representation conflicts.
Appendix A Limitations
While our method successfully reduces the tension between text and video embeddings by optimizing pair-wise $\Delta_{ij}$ based on the direction that minimizes the InfoNCE loss, several limitations remain:
A.1 Lack of Modality Alignment
The fundamental issue of modality gap persists. Despite reducing the tension between the embeddings of text and video, the two modalities still reside in completely disjoint regions of the representation space. Our method alleviates the modality gap by offloading some of the gradient tension, but it does not align the two modalities in a meaningful way. The solution we propose does not explicitly enforce a shared representation space between text and video; rather, it reduces the gradient tension through the optimization of $\Delta_{ij}$. As a result, the method mitigates the effects of the modality gap and reduces some noise, but it does not address the root cause of the misalignment between the two modalities.
A.2 Lack of Generalized Supervision for $\Delta$
Our approach relies heavily on adjusting $\Delta_{ij}$ through gradient-based optimization, but $\Delta_{ij}$ lacks a more generalizable supervision signal. Currently, we are optimizing each pair-wise $\Delta_{ij}$ based solely on the gradient of the InfoNCE loss, which only loosely guides the optimization direction. While this helps alleviate the tension between positive and negative pairs, it does not provide a stronger supervisory signal to guide the model toward a better generalization across unseen data.
In summary, while the method presents an effective way of mitigating gradient tension and improving the performance of contrastive learning, these limitations highlight areas for further development in future work.
Appendix B More Ablation Experiments
In the next four experiments conducted on the MSR-VTT-1KA dataset, we investigated the impact of different combinations of $c$ and $\eta$, the impact of the scale coefficient $\sigma$ in the directional diversity regularization, the impact of the lower bound of Norm-Based Regularization of Trust-Region Radii and the impact of the anchor choice of $\mathcal{L}_{\text{IB}}$.
B.1 Combination of Context Modality type $c$ and $\eta$ Choices
Table 4: Ablation on combinations of $c$ and $\eta$ (modality/context choice).
setting R@1↑ R@5↑ R@10↑ MdR↓ MnR↓
$c=t_{i}$ $\eta=-1$ 46.6 74.2 83.8 2.0 12.9
$c=t_{i}$ $\eta=+1$ 48.2 74.2 83.2 2.0 12.9
$c=v_{j}$ $\eta=-1$ 47.8 74.1 82.8 2.0 12.4
$c=v_{j}$ $\eta=+1$ 49.1 74.7 83.6 2.0 12.0
As shown in Table 4, the results show: 1) Fine-tuning the text side is better: By comparing different configurations of $c$ and $\eta$, fine-tuning the text side consistently outperforms other choices. Specifically, when $c=t_{i}$ and $\eta=1$, the text representation adjusts the increment $\Delta$ effectively, allowing for a higher semantic distinction of the increment. 2) Choosing $v$ as the context for $\psi$ is better: In this experiment, choosing the video modality $v_{j}$ as the context for $\psi$ yielded better results, especially when $c=v_{j}$ and $\eta=1$, achieving the highest R@1 score. This suggests that learning more fine-grained representations from the video modality helps improve cross-modal interactions and boosts performance.
Since videos in the MSR-VTT dataset are often clips from the same source video, leading to a high number of similar video samples, fine-grained learning becomes crucial. Adding this fine-grained adjustment to the text side enables $\Delta$ to have better semantic distinction and enhances retrieval performance.
B.2 Ablation of the Regularization Scale Coefficient $\sigma$ in Directional Diversity
As shown in Table 5, the results indicate: 1) Smaller $\sigma$ (0.5): When $\sigma$ is small, the diversity of the delta vectors is limited, which weakens the model’s ability to learn fine-grained representations. Although the R@1 score is 47.6, the model’s ability to distinguish between samples is reduced due to limited directional diversity. 2) Moderate $\sigma$ (1.0 and 2.0): In these configurations, the directional diversity of delta vectors is more balanced, effectively enhancing the model’s ability to capture cross-modal variations.
Table 5: Ablation on scale coefficient $\sigma$ for directional diversity.
$\sigma$ R@1↑ R@5↑ R@10↑ MdR↓ MnR↓
0.5 47.6 74.5 83.4 2.0 12.1
1.0 48.5 74.4 83.9 2.0 12.1
2.0 49.1 74.7 83.6 2.0 12.0
3.0 46.2 74.4 82.9 2.0 12.1
4.0 47.0 74.3 83.4 2.0 12.1
5.0 47.2 74.1 82.8 2.0 12.4
The R@1 score improves to 49.1, showing that appropriate variance increases retrieval performance. 3) Larger $\sigma$ (3.0 and 5.0): When $\sigma$ is too large, the directional difference between delta vectors becomes excessive, introducing noise and decreasing model performance. Specifically, with $\sigma=3.0$, the R@1 score drops to 46.2.
From this experiment, it is evident that an appropriately moderate $\sigma$ enables delta vectors to have useful directional diversity, thereby improving cross-modal representation learning. However, excessively large $\sigma$ may result in noise that negatively impacts model performance.
B.3 Ablation on $\Delta$ Norm Regularization Strength
Figure 3: R@1 score with varying $\lambda$ for $\Delta$ norm regularization.
We investigate the effect of varying the lower bound factor $\lambda$, which controls the target margin of $\varepsilon$ separation within each anchor $t_{i}$. As shown in Figure 3, increasing $\lambda$ from 0.1 to 0.5 improves R@1, with the best performance at $\lambda=0.5$. This suggests that moderate diversity in $\varepsilon_{ij}$ is beneficial for enhancing semantic discrimination. However, as $\lambda$ increases further, performance degrades, likely due to over-regularization.
To mitigate the instability introduced by hard truncation (i.e., directly thresholding $\varepsilon_{ij}$), we also experiment with a smooth approximation using a log-sum-exp formulation:
$\mathcal{L}_{\varepsilon-\text{LogSumExp}}=\log\left(1+\frac{1}{B}\sum_{j=1}^{B}\exp\left(-\left(\varepsilon_{ij}-\overline{\varepsilon}_{i}\right)^{2}\right)\right),\quad\overline{\varepsilon}_{i}=\frac{1}{B}\sum_{j=1}^{B}\varepsilon_{ij}$
Although this variant imposes a natural lower bound and provides vanishing gradients near convergence, it does not significantly improve retrieval. We hypothesize that this is due to gradient saturation when the variance among $\varepsilon_{ij}$ becomes too large, which in turn weakens the ability to further enforce directional separation. Unlike contrastive margins that encourage explicit pairwise gaps, the log-sum-exp loss lacks sufficient directional tension to align with the InfoNCE objective.
B.4 Effect of Anchor Choice in $\mathcal{L}_{\text{IB}}$
Table 6: Effect of anchor choice in KL regularization $\mathcal{L}_{\text{IB}}$. Anchoring on $v$ yields better performance.
Anchor R@1 R@5 R@10 MedR MnR
$t_{i}$ (text) 47.9 74.9 83.8 2.0 12.5
$v_{j}$ (video) 49.1 74.7 83.6 2.0 12.0
To examine the role of modality anchoring in the information bottleneck regularization, we compare two variants of the KL divergence loss $\mathcal{L}_{\text{IB}}$ applied on the learned increments $\Delta_{ij}$. In both cases, $\Delta_{ij}$ is modeled as a Gaussian variable, but we anchor its distribution either from the video side (i.e., conditionally modeling $\Delta$ given $v_{j}$) or the text side ($t_{i}$). Results are shown in Table 6.
We observe that using the video as the anchor yields significantly better retrieval performance. This suggests that the information bottleneck not only suppresses unnecessary variance in $\Delta$, but also guides the model to extract task-relevant signals from the noisier modality—videos in this case. Since $\Delta$ is generated based on the initial pairwise difference ($v_{j}-t_{i}$ or $t_{i}-v_{j}$), it inherently reflects the model’s prior on how to reconcile the cross-modal gap. Videos tend to carry more diverse and redundant visual information, and anchoring the KL constraint on $v_{j}$ allows the model to compress this input through $\Delta$, effectively learning a semantically meaningful adjustment. Conversely, anchoring on $t_{i}$ fails to suppress visual redundancy, and the resulting $\Delta$ captures less structured or discriminative transformations, thereby degrading alignment quality. This analysis highlights that $\Delta$ acts not only as a tension offloader and false-negative noise suppressor, but also as a structured bottleneck pathway between text and video.
Appendix C Analysis of $\Delta$ Behavior During Training
To better understand the role and dynamics of the learned modality-gap vector $\Delta$ throughout training, we visualize four key aspects of $\Delta$’s behavior, presented in Figure A1 to A4. These analyses reveal the underlying geometric transformations and provide further insight into how $\Delta$ facilitates optimization under InfoNCE loss. We will continue conducting ablation studies on the Text-to-Video retrieval task using the MSR-VTT [45] dataset.
C.1 Angle between $\Delta$ and $(v_{j}-t_{i})$
(a) Angle between $\Delta$ and $(v_{j}-t_{i})$
(b) Angle between $\Delta$ and $t_{i}$
(c) Norms of $t_{i}$ and $t_{\Delta_{ij}}$
(d) Pair distances: $t_{i}$/$t_{\Delta}$ vs. $v_{j}$
Figure 4: Training dynamics of the learned modality-gap vector $\Delta$. $t_{\text{new}}$ denotes $t_{\Delta}$. (a) $\Delta$ grows increasingly orthogonal to the initial modality gap, indicating embedding space expansion. (b) $\Delta$ initially aligns with the anchor $t_{i}$, then deviates to encode semantic distinctions. (c) Updated embeddings $t_{\text{new}}$ operate under larger norms, enlarging the contrastive space. (d) Pairwise distances increase and then stabilize, reflecting semantic separation and convergence.
As shown in Figure 4(a), we track the angle between $\Delta$ and the initial cross-modal gap vector $v_{j}-t_{i}$ throughout training, for both positive and negative pairs. At initialization, this angle is close to $\frac{\pi}{2}$, indicating that $\Delta$ is nearly orthogonal to $v_{j}-t_{i}$, and thus carries no meaningful alignment with the cross-modal semantic gap. This implies that the early $\Delta$ vectors do not differentiate between positive and negative pairs, acting more like isotropic perturbations in space. As training proceeds, however, this angle gradually increases into the obtuse region for both positive and negative samples. This reflects a significant shift in behavior—rather than attempting to directly bridge the modality gap $v_{j}-t_{i}$, the model learns to push $t_{i}$ away from $v_{j}$, effectively offloading the gradient tension induced by the modality gap and false negative interference. This offloading allows contrastive learning to take place in an expanded embedding space, where $\Delta$ modulates the representation geometry to ease the optimization burden. This trend is corroborated by Figure 4(d), where the Euclidean distances between the updated text embeddings $t_{\Delta_{ij}}=t_{i}+\Delta_{ij}$ and $v_{j}$ become significantly larger than the original distances $\|t_{i}-v_{j}\|$, for both positive and negative samples. That is, $\Delta$ introduces a global scaling effect in the representation space, and contrastive optimization is carried out under a larger geometric regime.
Interestingly, the new positive pair distances $\|t_{\Delta_{ii}}-v_{i}\|$ are also larger than their original counterparts $\|t_{i}-v_{i}\|$. Though this seems counterintuitive—since the gradient of InfoNCE with respect to $\Delta$ pushes toward $v_{i}$ — it actually reflects the relative nature of the InfoNCE loss. The network does not aim to minimize absolute distances, but rather to increase the similarity margin between matched and mismatched pairs. Thus, pushing all embeddings outward in norm (as further validated in Figure 4(c)) gives the model more room to maneuver in angular space, while the cosine similarity objective remains stable under such rescaling. This aligns with our design intent: to leverage $\Delta$ as a structural carrier for modality-aware tension redistribution, providing optimization flexibility on a normalized manifold.
Moreover, in early training stages, the angles between $\Delta$ and $v_{j}-t_{i}$ are nearly identical for positive and negative samples, showing that the model has not yet learned to encode fine-grained pairwise differences. But as training advances, a slight but consistent gap emerges—$\Delta$ for positive samples tends to have slightly smaller angles than that of negatives. This subtle divergence signals that $\Delta$ has begun to capture semantically meaningful pair-level distinctions, enabling more discriminative alignment in later stages.
C.2 Angle between $\Delta$ and $t_{i}$
As shown in Figure 4(b), the angle between $\Delta$ and $t_{i}$ exhibits a distinct pattern: it decreases rapidly in the early training phase, indicating that $\Delta$ is initially aligned with the anchor text embedding. This suggests that the model’s default behavior is to trust the prior structure of the text space and apply similar $\Delta$ directions across different $v_{j}$, especially when pairwise semantic differences are not yet learned.
However, over time, this alignment loosens—$\Delta$ deviates further from $t_{i}$ as the model begins to adapt to modality-specific differences. This marks a transition where the network $\psi$ no longer treats the anchor $t_{i}$ as a default direction and instead generates $\Delta$ based on nuanced distinctions across the visual features $v_{j}$, thus leading to more expressive and semantically grounded $\Delta$ vectors.
C.3 Pair Distances between Text and Video Features
Figure 4(d) visualizes the Euclidean distances between text and video pairs over the course of training, comparing both the original pairs $t_{i}\leftrightarrow v_{j}$ and the updated pairs $t_{\Delta_{ij}}\leftrightarrow v_{j}$, across positive and negative samples. Several important patterns emerge: First, we observe that positive pair distances remain consistently smaller than negative pair distances, both in the original and updated forms. This confirms that the learned $\Delta$ not only preserves the basic alignment structure of contrastive learning but also enhances semantic discrimination, effectively encoding fine-grained pairwise structure through the transformation. Second, for both positive and negative pairs, the distances between $t_{\Delta}$ and $v$ are larger than those between the original $t$ and $v$, demonstrating that the model pushes the updated text embeddings into a larger-scale embedding regime. This is precisely in line with our gradient tension offloading design: $\Delta$ introduces a controlled displacement that alleviates the direct optimization pressure on $t_{i}$, allowing contrastive comparisons to operate within a higher-norm, more expressive space, as also supported by Figure 4(c). The evolution of the distance curves over time also reflects meaningful training dynamics. Initially, all distances decrease, which we interpret as a normalization phase — embedding distributions at initialization are noisy and unstructured, and the model first compresses them into a tighter, more consistent geometric configuration. This is followed by a steady increase in distances, as the model begins to explicitly separate positive and negative samples to satisfy the InfoNCE objective. Finally, all curves converge and stabilize into a bounded range, suggesting that the semantic configuration of the embedding space has reached a relatively converged structural state.
Taken together, these observations highlight the critical role of $\Delta$ not only in scaling the embedding geometry but also in encoding structurally-aware, semantically-aligned displacements that facilitate tension redistribution and enable robust representation learning.
Appendix D Why We Don’t Use an Initial $\Delta_{ij}^{*}=0$ Expansion and the Need for a Non-Zero Initial $\Delta_{ij}^{*}$
D.1 Introduction to the Issue
In the previous approach, we considered using multiple variables $\Delta_{ij}$ to distribute the gradient impact of each video $v_{j}$ on the text embedding $t_{i}$. However, the specific form of $\Delta_{ij}$ was unclear. The idea was to update each $\Delta_{ij}$ in the direction that would reduce the InfoNCE loss. A natural first step was to expand the loss function $\mathcal{L}_{i}$ as a Taylor expansion around the initial state $\Delta_{ij}^{*}=0$, which would correspond to an expansion solely based on $t_{i}$.
However, this expansion introduces a key issue: the gradient terms in each $\Delta_{ij}$ expansion are computed assuming that all other $\Delta_{ik}$ are zero. This neglects the interactions between different $\Delta_{ij}$ and effectively decouples the updates for each pair. Since InfoNCE loss requires comparing relative similarities between different pairs, this decoupled approach violates the fundamental nature of the contrastive learning objective, as it does not respect the relative comparisons between the cosine similarities of different text-video pairs.
D.2 Why Expanding with $\Delta_{ij}^{*}=0$ Is Inadequate
When $\Delta_{ij}^{*}=0$ for all $j$, the gradient expansion for each pair $\Delta_{ij}$ looks like:
$\mathcal{L}_{i}(\Delta_{ij}^{*})\approx\mathcal{L}_{i}(0)+\sum_{j=1}^{B}\left[\nabla_{\!\Delta_{ij}}\mathcal{L}_{i}(0)\right]^{\top}\Delta_{ij},$ (17)
where $\nabla_{\!\Delta_{ij}}\mathcal{L}_{i}(0)$ is the gradient of the loss evaluated at $\Delta_{ij}=0$. In this case, the gradient $p_{ij}-y_{ij}$ is computed as:
$\nabla_{\!\Delta_{ij}}\mathcal{L}_{i}(0)=\left(p_{ij}(0)-y_{ij}\right)\nabla_{\!\Delta_{ij}}\cos_{ij}(0),$ (18)
where $p_{ij}(0)$ is the softmax term evaluated at the cosine similarity $\cos(t_{i},v_{j})$. This expansion does not account for the interdependence between different $\Delta_{ik}$, and the gradients are computed as if each $\Delta_{ij}$ were independent of all other $\Delta_{ik}$, which violates the relative comparison required for contrastive learning.
D.3 The Need for a Non-Zero Initial $\Delta_{ij}^{*}$
To avoid this issue, we recognize that each $\Delta_{ij}$ should be updated in a way that reflects the interdependence between different pairs, not just based on an independent text-video pair. The gradient of the InfoNCE loss must be evaluated with respect to the current state of all $\Delta_{ij}$, not just the initial zero state.
Thus, we need to initialize $\Delta_{ij}^{*}$ in a way that reflects the current state of all other pairs rather than assuming they are zero. The update for each $\Delta_{ij}$ should respect the relationship between all the pairwise similarities, as the optimization of one pair affects the relative similarity between all other pairs.
In this case, the correct approach is to perform the first-order Taylor expansion around a non-zero state $\Delta_{ij}^{(t)}$ that has already been optimized for several steps. This allows us to incorporate the effect of each pair $\Delta_{ij}$ on the entire set of pairwise comparisons, thus preserving the relative relationships that are critical for InfoNCE loss.
D.4 Optimization Objective and Trust-Region Constrained Solution
Having established the need for a coupled, non-zero initialization of $\Delta_{ij}$ to preserve the relative structure required by InfoNCE, we now formulate the optimization problem for finding the optimal update direction under a trust region constraint. Specifically, we assume a trust region radius $\varepsilon$ that bounds the magnitude of each $\Delta_{ij}$ and analyze the solutions in both the single-step (non-iterative) and iterative update settings.
Non-Iterative First-Order Update
We consider the first-order Taylor expansion of the per-anchor loss $\mathcal{L}_{i}$ with respect to $\Delta_{ij}$ evaluated at the origin:
$\mathcal{L}_{i}(\Delta_{ij})\approx\mathcal{L}_{i}(0)+\nabla_{\Delta_{ij}}\mathcal{L}_{i}(0)^{\top}\cdot\Delta_{ij},$ (19)
subject to the trust-region constraint $\|\Delta_{ij}\|\leq\varepsilon$. This forms a standard constrained linear minimization problem, where the optimal $\Delta_{ij}$ minimizes the inner product with the gradient. By the Cauchy–Schwarz inequality:
$\nabla_{\!\Delta_{ij}}\mathcal{L}_{i}^{\top}\cdot\Delta_{ij}\geq-\|\nabla_{\!\Delta_{ij}}\mathcal{L}_{i}\|\cdot\|\Delta_{ij}\|,$ (20)
equality holds when $\Delta_{ij}$ and $-\nabla_{\Delta_{ij}}\mathcal{L}_{i}$ are colinear. Thus, the optimal update is in the steepest descent direction:
$\Delta_{ij}^{*}=-\varepsilon\cdot\frac{\nabla_{\Delta_{ij}}\mathcal{L}_{i}}{\|\nabla_{\Delta_{ij}}\mathcal{L}_{i}\|}.$ (21)
This solution, however, ignores the coupling between different $\Delta_{ik}$ and breaks the softmax structure, violating the relative comparison essential to InfoNCE, as discussed in the subsection above. Therefore, a more faithful solution must incorporate the influence of all $\Delta_{ij}$ jointly.
Iterative Update with Coupled Expansion
We instead consider the Taylor expansion centered at the current state $\Delta_{ij}^{(t)}$, capturing the coupled contributions of all other $\Delta_{ik}\neq\Delta_{ij}$. We seek the next update step:
$\Delta_{ij}^{(t+1)}=\Delta_{ij}^{(t)}+\delta,\quad\text{with }\|\Delta_{ij}^{(t+1)}\|\leq\varepsilon.$ (22)
We again minimize the first-order approximation:
$\mathcal{L}_{i}(\Delta_{ij}^{(t+1)})\approx\mathcal{L}_{i}(\Delta_{ij}^{(t)})+\nabla_{\!\Delta_{ij}}\mathcal{L}_{i}^{(t)}{}^{\top}\cdot(\Delta_{ij}^{(t+1)}-\Delta_{ij}^{(t)}).$
Letting $\hat{g}_{ij}^{(t)}=\nabla_{\Delta_{ij}}\mathcal{L}_{i}^{(t)}/\|\nabla_{\Delta_{ij}}\mathcal{L}_{i}^{(t)}\|$, we propose a step in the steepest descent direction:
$\Delta_{ij}^{(t+1)}=\Delta_{ij}^{(t)}-\alpha_{ij}^{(t)}\cdot\hat{g}_{ij}^{(t)}.$ (23)
To satisfy the norm constraint $\|\Delta_{ij}^{(t+1)}\|\leq\varepsilon$, we solve for the maximum allowable step size $\alpha_{ij}^{(t)}$ using the triangle and Cauchy–Schwarz inequalities:
$\|\Delta_{ij}^{(t)}-\alpha_{ij}^{(t)}\hat{g}_{ij}^{(t)}\|^{2}\leq\varepsilon^{2}.$ (24)
Expanding this yields the quadratic inequality:
$\alpha_{ij}^{(t)2}-2\alpha_{ij}^{(t)}\Delta_{ij}^{(t)\top}\hat{g}_{ij}^{(t)}+\|\Delta_{ij}^{(t)}\|^{2}-\varepsilon^{2}\leq 0,$ (25)
whose maximal feasible solution is:
$\alpha_{ij}^{(t)}=\Delta_{ij}^{(t)\top}\hat{g}_{ij}^{(t)}+\sqrt{\left(\Delta_{ij}^{(t)\top}\hat{g}_{ij}^{(t)}\right)^{2}-\|\Delta_{ij}^{(t)}\|^{2}+\varepsilon^{2}}.$ (26)
Thus, the update maintains the trust region while moving in the steepest descent direction under the coupled loss landscape.
D.5 Why Use Neural Networks for Approximation?
The reason we opt for neural networks to approximate the current state of $\Delta_{ij}$ rather than directly solving for each $\Delta_{ij}$ through numerical optimization is primarily due to efficiency and scalability:
Handling Non-Linearities: The relationship between the embeddings of text and video is highly non-linear. Neural networks are particularly well-suited for learning such complex, non-linear mappings, allowing the model to effectively learn the best updates for $\Delta_{ij}$ without requiring explicit mathematical formulations for each pairwise interaction.
Scalability: As the size of the dataset increases, the computational cost of directly solving each $\Delta_{ij}$ grows rapidly. Neural networks allow for scalable learning since once trained, they can quickly compute updates for large batches without re-solving the optimization problem for each individual pair, making them a more practical choice in real-world settings.
In short, neural networks allow for faster convergence and better scalability compared to direct numerical solutions, making them essential for large-scale contrastive learning tasks.
Appendix E Computational Complexity Analysis.
In our proposed GARE framework, each mini-batch of size $B$ involves the prediction of a pair-specific increment $\Delta_{ij}\in\mathbb{R}^{D}$ for all text-video pairs. This yields a forward computational complexity of $\mathcal{O}(B^{2}D)$, since the learned increment network $\psi$ generates $B\times B$ outputs of dimension $D$. While quadratic in batch size, this cost is efficiently amortized during training and can be parallelized on modern hardware.
Importantly, this design avoids the substantial overhead required by a numerical optimization scheme. If one were to directly minimize the first-order Taylor expansion of the loss with respect to $\Delta$ using explicit iterative updates, two key obstacles arise: (1) The expansion is strictly first-order and lacks nonlinear modeling capacity, making it poorly suited for generalizing to unseen test-time structures; (2) Due to the softmax-coupled nature of InfoNCE, the gradients of each $\Delta_{ij}$ are entangled with all other $\Delta_{ik}$, requiring $T$ rounds of joint optimization for a converged update. As a result, the overall complexity escalates to $\mathcal{O}(TB^{2}D)$, where $T$ denotes the number of inner optimization steps. In addition to the increased cost, this setup suffers from severe gradient coupling and lacks a principled way to impose per-pair trust region constraints such as $|\Delta_{ij}|\leq\varepsilon$.
By contrast, our network-based approach amortizes this optimization and learns to implicitly enforce structure-aware constraints via training, making it not only more efficient but also more generalizable.